OpenStudy (anonymous):

a leaky valve on the water meter over charges residents for one gallon of water every 1 1/3 months. the over charged amount w varies directly with time t. find the equation that models this direct variation. how many months it will take for the residents to be over charged for 5 gallons of water? A w=3/4t; 6 2/3 months B w= 4/3t; 6 2/3 months C w= 3/4t; 3 3/4 months D w= 4/3t; 3 3/4 months

ganeshie8 (ganeshie8):

4/3 months ---> 1 gallon w = 4/3 t -----------------------(1) 5 = 4/3 t solve t